# Generalization performance of narrow one-hidden layer networks in the teacher-student setting

### v3

@@ Abstract
Understanding the generalization abilities of neural networks for simple input-output distributions is crucial to account for their learning performance on real datasets. The classical teacher-student setting, where a network is trained from data obtained thanks to a label-generating teacher model, serves as a perfect theoretical test bed. In this context, a complete theoretical account of the performance of fully connected one-hidden layer networks in the presence of generic activation functions is lacking. In this work, we develop such a general theory for narrow networks, i.e. with a large number of hidden units, yet much smaller than the input dimension. Using methods from statistical physics, we provide closed-form expressions for the typical performance of both finite temperature (Bayesian) and empirical risk minimization estimators, in terms of a small number of summary statistics. In doing so, we highlight the presence of a transition where hidden neurons specialize when the number of samples is sufficiently large and proportional to the number of parameters of the network. Our theory accurately predicts the generalization error of neural networks trained on regression or classification tasks with either noisy full-batch gradient descent (Langevin dynamics) or full-batch gradient descent.
@@ 1 Introduction and related works
Predicting theoretically the generalization abilities of complex neural architectures for generic tasks in terms of the number of parameters and training samples is a daunting task. The statistical physics approach has been successful in doing so in the past few decades [seung1992statistical, engel2001statistical, zdeborova2020understanding], where the performance of a network is characterized by the typical generalization error over synthetic data generated by probabilistic models of increasing statistical complexity. The replica method from the physics of disordered systems [mezard1987spin, charbonneau2023spin] is particularly effective for studying feed-forward networks solving supervised learning tasks in the high-dimensional regime. The two classical, complementary settings in this area are the storage capacity problem [gardner_optimal_1988], where a network is tasked with memorizing a random input-output mapping, and the teacher-student scenario [schwarze_statistical_1992], where a rule mapping labels to input data is instantiated also in a form of a similar random network architecture.
@@
While the analysis of the typical performance of single-layer networks (perceptrons) is – at least in the case of continuous parameters – relatively straightforward and classic [gardner_space_1988], the one-hidden layer case is much more challenging. In particular, research has been mainly focused on the case of committee machines, where the second-layer weights are not trained. A variety of works (detailed in the next section) analyzed the information-theoretic limits of supervised learning in committee machines, where a network is trained on $P$ independent $N$-dimensional examples, with the sample complexity parameter $\alpha:=P/N=O_{N}(1)$ and $N\to\infty$. The general picture emerging is that generalization performance is controlled by the ratio of the dataset size and the number of network parameters. In particular, specialization of hidden neurons in the direction of the ones in the teacher can only occur when the number of samples is of the order of the number of network parameters, i.e. $P\propto NK$.
@@
Classical results exist for the simpler case of tree architectures (with non-overlapping receptive fields of the hidden units) and for fully connected (FC) committees with sign activation function [schwarze_statistical_1992, barkai_broken_1992, schwarze_learning_1993, schwarze_generalization_1992]. While recent studies [oostwal_hidden_2021, citton_phase_2025] started addressing the case of arbitrary activations, a general theory of learning using empirical risk minimization (ERM) is not yet present in the statistical mechanics literature. In this work, we study the learning problem in the limit where the input dimension $N$ diverge and the number of hidden units $K$ is large, yet much smaller than $N$. As we will show, such $1\ll K\ll N$ regime lends itself to a typical-case analysis using the replica method, in terms of a finite number of order parameters (i.e., sufficient statistics to describe the generalisation error). We therefore study the typical performance of ERM estimators in a teacher-student setup with generic activation function, investigating the emergence of learning phase transitions. We obtain closed form expressions for the weight statistics and the overlap between the teacher and student weight vectors, thus providing a theoretical prediction for the generalization error as a function of sample complexity, expressed as the ratio of training set size and the total number of adjustable weights in the network. We show that our theory predicts well the performance of the Langevin dynamic (LD), including in the limit of vanishing noise, but also the test-time behavior even when networks are trained with pure gradient descent (GD). Our analysis reveals distinct regimes of learning, including a computationally hard phase (or “statistical-to-computational gap”) in which posterior sampling methods such as LD converge extremely slowly. To probe this regime, we also study an informed initialization of LD, where the student weights are initialized close to the teacher configuration, allowing the dynamics to access specialized states that are otherwise difficult to sample.
@@
Shallow neural networks have been extensively studied using a statistical physics framework [Cui_narrow_2025], where a description of the typical learning and generalization behavior over a joint input-label ensemble can be provided in the limit of large input dimension $N\to\infty$. The specific phenomenology of such a problem is related to the number of hidden units.
@@
In the $K=O_{N}(1)$ regime, a two-layer neural network falls into the category of multi-index functions. Classical works studied online learning dynamics in such a model, deriving a set of ordinary differential equations for the sufficient statistics of the hidden representations [saad_solla_online, saad_solla_solution, Riegler_online, Biehl_online, goldt_dynamics]. In the teacher-student setting, where all information about the teacher is available to the student except for the input-to-hidden-layer weights, a rigorous classification of the difficulty in recovering the target function via approximate message passing (AMP)/TAP-like algorithms has been established in [troiani_fundamental_2024]. The work [abbe_sgd_2023] systematically classified SGD performance on two-layer neural networks for any class of multi-index functions based on the leap complexity of the target function.
@@
While many questions have been resolved for finite index, the large-index $K$ regime remains to be explored. The regime $1\ll K\ll N$ has been addressed both in the tree [monasson_learning_1995, sollich_learning_1996, urbanczik1994tree] and the FC case [schwarze_statistical_1992, schwarze_learning_1993, engel_storage_1992, schwarze_discontinuous_1993, schwarze_learning_1993-1, urbanczik1997full]. These classical studies focused on sign activation function (or erf  [ahr1999statistical]) activation – with either binary or spherically constrained first-layer weights – and i.i.d. inputs. In the FC case, [Ahr_committee_1999] studiedthe effect of $L_{2}$ regularization using an annealed approximation, corresponding to the large-dataset limit. A FC model with a bounded $\mathcal{C}^{2}$ activation function with bounded first and second derivatives was studied from a Bayesian-optimal perspective in [aubin_committee_2018], where the authors supported replica predictions with rigorous theorems, proposing an efficient inference algorithm based on AMP. In the ERM context, a simple kernel order parameter has been shown to control the storage capacity in the simplified case of a tree committee with generic activation functions [baldassi_properties_2019, zavathone_activation, annesi_fullrsb]. Using similar methods, a recent work identified analogous kernel order parameters for generic activation function and a particular choice of second-layer weights [nishiyama_solution_2025]. We further note the works [oostwal_hidden_2021, citton_phase_2025], which analyzed networks with square-summable activation functions using Hermite expansions and a simplifying approximation in the limit of a very large random dataset. We also mention a recent burst of interest for the extensive-width regime $K=O(N)$ [maillard_quadratic, barbier2025statistical, montanari2025dynamical, ren2025emergence]. In particular, when comparing our work with [barbier2025statistical], a key distinction in the extensive-$K$ regime is the emergence of kernel-like learning in the small-data regime (which does not occur in our small-to-medium-$K$ setting), as opposed to specialization in the high-$\alpha$ regime, which appears in both scalings.
@@
Our contributions — In this work we develop a complete theoretical account of the typical performance of a one-hidden-layer neural network in the regime $1\ll K\ll N$ in the presence of i.i.d. inputs and responses generated from a teacher network with the same structure. Our contributions are the following:
@@
  • We compute the generalization error and the corresponding learning curves as a function of the number of samples, for committee machines with a generic activation function and a generic loss, in the classic teacher-student setting. Unlike previous works, which used an annealed calculation [oostwal_hidden_2021, citton_phase_2025] for regression, our approach relies on a quenched computation of the free entropy using the replica method. This computation is of independent interest to the machine learning community, and we anticipate that it will have further applications in learning theory.
  • We validate our theoretical predictions in the high-dimensional limit through Langevin-based simulations.
@@
  • We empirically compare the performance of ERM under LD and GD. While LD can sample from the posterior when run long enough, GD is a deterministic optimizer and does not explore the full posterior. Remarkably, we find that GD converges to solutions whose generalization performance is equivalent to those obtained by LD, suggesting that optimization and sampling dynamics can yield statistically similar learned representations.
Finally, we summarize these regimes in a novel phase diagram that captures specialization transitions and the effect of regularization across different sample complexities.
@@ 2.1 Empirical risk minimization, and statistical physics formulation
Throughout the paper, we consider the standard supervised learning setup with a synthetic dataset $\mathcal{D}:=\{(\mathbf{x}^{\mu},y^{\mu}_{\star})\}_{\mu=1}^{P}$, with responses generated by a two-layer teacher neural network. The samples are thus generated as follows: (i) Construct the target function by sampling entry-wise i.i.d. the first-layer teacher weights $\mathbf{W}^{\star}=(\mathbf{w}_{k}^{\star}\in\mathbb{R}^{N})_{k=1}^{K}\in\mathbb{R}^{K\times N}$ from $P_{W^{\star}}$, and the second-layer weights $\boldsymbol{A}^{\star}=(A_{k}^{\star})_{k=1}^{K}$ from $P_{A}$, with finite first and second moments. (ii) The inputs $\mathbf{x}^{\mu}$ are i.i.d. standard Gaussian vectors: $\mathbf{x}^{\mu}\sim\mathcal{N}(0,\mathbf{I}_{N})$. (iii) The responses/labels $(y^{\mu}_{\star})_{\mu=1}^{P}$ are generated as
@@
The starting point of the statistical physics approach to learning is to compute the log-partition function, the so-called free entropy: $\ln\mathcal{Z}_{\beta}/(NK)$. We expect it to be self-averaging, i.e. its variance w.r.t. the realization of the problem (teacher and data) vanishes as $N\to\infty$ (this is proven using standard concentration techniques, see e.g. [barbier2019optimal]). This allows us to focus on its expectation value
@@ 2.2 Main results: closed form formulas for the free entropy, test and training errors
where $m$ denotes a $K\times K$ matrix whose elements can be expressed in terms of two parameters $m_{d}$ and $m_{a}$ as $m=m_{d}\mathbf{I}_{K}+(m_{a}/K)\mathbf{1}_{K}\mathbf{1}_{K}^{\intercal}$ and similarly for $v,q$. The “entropic potential” is
@@
where ${\mathbf{W}}_{a}$ and ${\mathbf{W}}_{b}$ are conditionally (on $\mathcal{D}$) i.i.d. samples from the Gibbs measure. The sufficient statistics are thus three $K\times K$ matrices whose physical meaning are: $q^{*}$, the student-student overlap, quantifies the average correlation between two different weight samples from the Gibbs measure (5); $q_{\text{self}}^{*}$, the self-student overlap, measures the correlation matrix between learned weight vectors of a single configuration; and finally $m^{*}$, the teacher-student overlap, describes the alignment between the learned weights and the teacher’s. In the high-dimensional limit, our theory assumes that these matrices have constant diagonal and off-diagonal entries (up to irrelevant fluctuations), with the latter scaling differently with $K$. This assumption, along with the application of the central limit theorem following the approach in tree-like committee machines [baldassi_properties_2019, zavathone_activation], leads to a substantial simplification of the RS entropy and enables the treatment of general activation functions for large $K$, see the supplementary material (SM) for the complete derivation.
@@
Given $(m^{*},q^{*},v^{*})$, a solution of the RS equations maximizing the RS potential (9), the RS approximation to the (Gibbs) generalisation error $\lim_{N\to\infty}\epsilon_{\rm Gibbs}$ is given by $\mathbb{E}_{(t,s)}[(f(t)-f(s))^{2}]$. Notice that $\mathcal{K}_{0}(d_{1},d_{2},a)$ coincides with the NNGP kernel [Neal1996, Williams1996] of large-width neural networks at initialization. Similarly, the RS approximation to the training error $\lim_{N\to\infty}\epsilon_{\rm tr}$ reads $\partial_{\beta}\mathcal{G}_{E}|_{m^{*},q^{*},v^{*}}$.
Figure 1: Free entropy (first panel) and order parameters associated with the stationary solutions of the RS equations, as a function of the sample complexity $\tilde{\alpha}$, for quadratic activation function $\sigma(x)=x^{2}$, MSE loss, weight decay $\lambda=0.1$, number of hidden units $K=10$ and $\beta=10$. Blue curves represent off-diagonal components of the teacher-student overlap $m^{*}$, the student-student overlap $q^{*}$, and the self-student overlap $q_{\rm self}^{*}$ Red curves represent the dominant diagonal components of these quantities. The dotted-dashed lines correspond to the permutation-symmetric branch (where $m_{d}^{*}=q_{d}^{*}=0$), the solid lines to the specialization branch ($m_{d}^{*}>0,q_{d}^{*}>0$) and the dashed one to the memorization branch ($m_{d}^{*}=0,q_{d}^{*}>0$).

### v4

@@ Abstract
Understanding the generalization properties of neural networks on simple input–output distributions is key to explaining their performance on real datasets. The classical teacher–student setting, where a network is trained on data generated by a teacher model, provides a canonical theoretical test bed. In this context, a complete theoretical characterization of fully connected one-hidden-layer networks with generic activation functions remains missing. In this work, we develop a general framework for such networks with large width, yet much smaller than the input dimension. Using methods from statistical physics, we derive closed-form expressions for the typical performance of both finite-temperature (Bayesian) and empirical risk minimization estimators in terms of a small number of order parameters. We uncover a transition to a specialization phase, where hidden neurons align with teacher features once the number of samples becomes sufficiently large and proportional to the number of network parameters. Our theory accurately predicts the generalization error of networks trained on regression and classification tasks using either noisy full-batch gradient descent (Langevin dynamics) or deterministic full-batch gradient descent.
@@ 1 Introduction and related works
Predicting theoretically the generalization abilities of complex neural architectures for generic tasks in terms of the number of parameters and training samples is a daunting task. The statistical physics approach has been successful in doing so in the past few decades [46, 16, 54], where the performance of a network is characterized by the typical generalization error over synthetic data generated by probabilistic models of increasing statistical complexity. The replica method from the physics of disordered systems [27, 12] is particularly effective for studying feed-forward networks solving supervised learning tasks in the high-dimensional regime. The two classical, complementary settings in this area are the storage capacity problem [17], where a network is tasked with memorizing a random input-output mapping, and the teacher-student scenario [44], where a rule mapping labels to input data is instantiated also in a form of a similar random network architecture.
@@
While the analysis of the typical performance of single-layer networks (perceptrons) is – at least in the case of continuous parameters – relatively straightforward and classic [18], the one-hidden layer case is much more challenging. In particular, research has been mainly focused on the case of committee machines, where the second-layer weights are not trained. A variety of works (detailed in the next section) analyzed the information-theoretic limits of supervised learning in committee machines, where a network is trained on $P$ independent $N$-dimensional examples, with the sample-to-parameter ratio $\alpha:=P/N=O_{N}(1)$ and $N\to\infty$ . The general picture emerging is that generalization performance is controlled by the ratio of the dataset size and the number of network parameters. In particular, specialization of hidden neurons in the direction of the ones in the teacher can only occur when the number of samples is of the order of the number of network parameters, i.e. $P\propto NK$.
Classical results exist for the simpler case of tree architectures (with non-overlapping receptive fields of the hidden units) and for fully connected (FC) committees with sign activation function [44, 10, 43, 41]. While recent studies [35, 13] started addressing the case of arbitrary activations, a general theory of learning using empirical risk minimization (ERM) is not yet present in the statistical mechanics literature. This gap is particularly relevant for modeling more realistic machine-learning tasks, where the learner lacks access to the true data-generating distribution—unlike in Bayes-optimal settings. This framework allows for the exploration of mismatched priors and provides a pathway to understand how emergent neural representations adapt to the underlying structure of the data. In this work, we study the learning problem in the limit where the input dimension $N$ diverge and the number of hidden units $K$ is large, yet much smaller than $N$. More precisely, we consider the regime $1\ll K\ll N$, taking $N\to\infty$ first and subsequently analyzing the large-$K$ limit. While this asymptotic setting formally allows for joint scalings such as $K=o(N)$, our analysis is controlled in the sequential limit described above, and the joint-scaling regime should be understood as a consistent asymptotic extension. This scaling enables a typical-case analysis via the replica method in terms of a finite number of order parameters (i.e., sufficient statistics that characterize the generalization error).
@@
We therefore study the typical performance of ERM estimators in a teacher-student setup with generic activation function and fixed (non-trainable) second-layer weights, investigating the emergence of learning phase transitions. We obtain closed form expressions for the weight statistics and the overlap between the teacher and student weight vectors, thus providing a theoretical prediction for the generalization error as a function of sample-to-parameter ratio $\tilde{\alpha}:=\alpha/K$, expressed as the ratio of training set size and the total number of adjustable weights in the network. We show that our framework accurately predicts the performance of Langevin dynamics (LD), including in the vanishing-noise limit, and also captures the test-time behavior of networks trained with pure gradient descent (GD). Our analysis identifies distinct learning regimes, notably a computationally hard phase (i.e., a statistical–computational gap) in which posterior sampling methods such as LD exhibit exponentially slow convergence. To probe this regime, we consider an informed initialization of LD, where the student weights are initialized in the vicinity of the teacher configuration, enabling the dynamics to reach specialized states that are otherwise algorithmically hard to sample.
@@
Shallow neural networks have been extensively studied using a statistical physics framework [14], where a description of the typical learning and generalization behavior over a joint input-label ensemble can be provided in the limit of large input dimension $N\to\infty$. The specific phenomenology of such a problem is related to the number of hidden units.
@@
In the $K=O_{N}(1)$ regime, a two-layer neural network falls into the category of multi-index functions. Classical works studied online learning dynamics in such a model, deriving a set of ordinary differential equations for the sufficient statistics of the hidden representations [40, 39, 38, 11, 19]. In the teacher-student setting, where all information about the teacher is available to the student except for the input-to-hidden-layer weights, a rigorous classification of the difficulty in recovering the target function via approximate message passing (AMP)/TAP-like algorithms has been established in [49]. The work [1] systematically classified SGD performance on two-layer neural networks for any class of multi-index functions based on the leap complexity of the target function.
@@
While many questions have been resolved for finite index, the large-index $K$ regime remains to be explored. The regime $1\ll K\ll N$ has been addressed both in the tree [29, 48, 50] and the FC case [44, 43, 15, 45, 42, 51]. These classical studies focused on sign (or erf  [3]) activation function – with either binary or spherically constrained first-layer weights – and i.i.d. inputs. In the FC case, [2] studied the effect of $L_{2}$ regularization using an annealed approximation, corresponding to the large-dataset limit. A FC model with a bounded $\mathcal{C}^{2}$ activation function with bounded first and second derivatives was studied from a Bayesian-optimal perspective in [6], where the authors supported replica predictions with rigorous theorems, proposing an efficient inference algorithm based on AMP. In the ERM context, a simple kernel order parameter has been shown to control the storage capacity in the simplified case of a tree committee with generic activation functions [7, 53, 4]. Using similar methods, a recent work identified analogous kernel order parameters for generic activation function and a particular choice of second-layer weights [34]. We further note the works [35, 13], which analyzed networks with square-summable activation functions using Hermite expansions and a simplifying approximation in the limit of a very large random dataset. Finally, we also mention a recent burst of interest for the extensive-width regime $K=O(N)$ [25, 8, 31, 37]. In particular, when comparing our work with [8], a key distinction in the extensive-$K$ regime is the emergence of kernel-like learning in the small-data regime (which does not occur in our narrow-but-large-$K$ setting). Conversely, both scalings converge on specialization in the high-$\alpha$ regime. We emphasize that, based on our theoretical observations, for large sample-to-parameter ratio, the Bayes-optimal generalization error (BOE) in our framework is consistent with the extensive width result of [8]. However, exact analytical agreement for all values of $\alpha$ requires $K/N\to 0$. In other cases, while the BOE may match, the underlying order parameters differ. Our approach remains analytically tractable by reducing the problem to a fixed set of scalar order parameters. This avoids the complex functional matrix order parameters– and the necessary integration of replica theory with random matrix models, specifically the Harish Chandra–Itzykson–Zuber (HCIZ) “spherical” integral [22, 20] – required in the $K=O(N)$ case.
@@
Our contributions — In this work, we develop a complete theoretical characterization of the typical performance of a one-hidden-layer neural network in the regime $1\ll K\ll N$ –again formally interpreted as a sequential limit $N$ large first then $K$ large– in the presence of i.i.d. inputs and responses generated from a teacher network with the same structure.
Our contributions are the following:
@@
  • We compute the generalization error and the corresponding learning curves as a function of the number of samples, for narrow and shallow networks with generic activation functions and a generic loss, in the classic teacher-student setting. Unlike previous works, which used an annealed calculation [35, 13] for regression, our approach relies on a quenched computation of the free entropy using the replica method, whose predictions have been validated through Langevin-based simulations. While our results agree with the annealed computation for very large dataset regime, finite datasets can display a continuous-to-discontinous behavior, as we report for the ReLU activation function. This computation is of independent interest to the machine learning community, and we anticipate that it will have further applications in learning theory.
@@
  • We study the limit in which the Bayesian posterior concentrates on minima of the empirical loss, and experimentally compare its predictions to solutions found by ERM using GD and LD. While LD can sample from the posterior when run long enough, GD is a deterministic optimizer and does not explore the full posterior. Remarkably, we find that GD converges to solutions whose generalization performance is equivalent to those obtained by LD, suggesting that optimization and sampling dynamics can yield statistically similar learned representations.
  • We provide a phase diagram highlighting the role of the data-to-parameters ratio $\tilde{\alpha}$ and the $L_{2}$ regularization strength $\lambda$ in shaping the properties of the equilibrium solution. This characterization identifies the optimal regularization required to approach the Bayes-optimal error.
  • In the Bayes-optimal setting, taking the $K\to\infty$ limit after $N\to\infty$ recovers the BOE reported in  [8] for activation functions without second Hermite polynomial in their Hermite decomposition. This suggests that for such activations, the large-width ($K$) and large-dimension ($N$) limits commute. For a general activation function, while both theories predict asymptotically (as $\tilde{\alpha}\to\infty$) the same generalization error, our results exactly match those of  [8] at all $\tilde{\alpha}$ only when $\gamma:=K/N\to 0$.
@@ 2.1 Empirical risk minimization, and statistical physics formulation
Throughout the paper, we consider the standard supervised learning setup with a synthetic dataset $\mathcal{D}:=\{(\mathbf{x}^{\mu},y^{\mu}_{\star})\}_{\mu=1}^{P}$, with responses generated by a two-layer teacher neural network. The samples are thus generated as follows: (i) Construct the target function by sampling entry-wise i.i.d. the first-layer teacher weights ${\mathbf{W}}^{\star}=(\mathbf{w}_{k}^{\star}\in\mathbb{R}^{N})_{k=1}^{K}\in\mathbb{R}^{K\times N}$ from $P_{W^{\star}}$, and the second-layer weights $\boldsymbol{A}^{\star}=(A_{k}^{\star})_{k=1}^{K}$ from $P_{A^{\star}}$, with finite first and second moments. (ii) The inputs $\mathbf{x}^{\mu}$ are i.i.d. standard Gaussian vectors: $\mathbf{x}^{\mu}\sim\mathcal{N}(0,\mathbf{I}_{N})$. (iii) The responses/labels $(y^{\mu}_{\star})_{\mu=1}^{P}$ are generated as
@@
The starting point of the statistical physics approach to learning is to compute the log-partition function, the so-called free entropy: $\ln\mathcal{Z}_{\beta}/(NK)$. We expect it to be self-averaging, i.e. its variance w.r.t. the realization of the problem (teacher and data) vanishes as $N\to\infty$ (this is proven using standard concentration techniques, see e.g. [9]). This allows us to focus on its expectation value
@@
The asymptotic regime in Eq. (4) formally allows for joint scaling of the network width with the input dimension, such as $K\propto N^{\delta}$ with $\delta<1$. However, the replica computation presented below is controlled in the sequential limit, where $N,P\to\infty$ a fix (but arbitrary) $K$. In this limit, the free entropy is obtained from an extremization over $K\times K$ overlap matrices. In a second step, we consider the regime of large but finite $K$. Here, we assume the overlap matrices are parameterized by a uniform diagonal and off-diagonal structure, as suggested in [41] and detailed in the following section. Upon rescaling the free entropy by $K$, the stationary equations admit a well-defined large-$K$ limit, yielding asymptotically equivalent results. Consequently, the joint-scaling regime ($K\propto N^{\delta}$ with $\delta<1$) should be interpreted a heuristic extension consistent with these asymptotic equations, rather than a rigorously derived limit within the current framework.
@@ 2.2 Main results: closed form formulas for the free entropy, test and training errors
where $m$ denotes a $K\times K$ matrix whose elements can be expressed in terms of two parameters $m_{d}$ and $m_{a}$ as $m=m_{d}\mathbf{I}_{K}+(m_{a}/K)\mathbf{1}_{K}\mathbf{1}_{K}^{\intercal}$ and similarly for $v,q$. This parametrization of the overlap matrices can be justified in the regime $K\ll N$. By invoking statistical permutation symmetry – naturally suggested by the indistinguishable nature of the hidden units – we expect overlaps between distinct units to concentrate. This allows us to describe the system using a reduced set of scalar order parameters rather than tracking every individual pair. This choice is further constrained by asymptotic self-consistency: to ensure the network’s output variance remain $\mathcal{O}(1)$ in the large $K$ limit, these off-diagonal overlaps must scales as $\mathcal{O}(1/K)$. If they were $\mathcal{O}(1)$, the K(K1)K(K-1) correlations between different units sum would cause a divergence. Thus, the $1/K$ scaling represents a structurally stable state where units are collectively balanced, and where individual fluctuations around their mean become negligible in the thermodynamic limit.
Then, the “entropic potential” is
@@
where ${{\mathbf{W}}}_{a}$ and ${{\mathbf{W}}}_{b}$ are conditionally (on $\mathcal{D}$) i.i.d. samples from the Gibbs measure. The sufficient statistics are thus three $K\times K$ matrices whose physical meaning are: $q^{*}$, the student-student overlap, quantifies the average correlation between two different weight samples from the Gibbs measure (5); $q_{\text{self}}^{*}$, the self-student overlap, measures the correlation matrix between learned weight vectors of a single configuration; and finally $m^{*}$, the teacher-student overlap, describes the alignment between the learned weights and the teacher’s. In the high-dimensional limit, our theory assumes that these matrices have constant diagonal and off-diagonal entries (up to irrelevant fluctuations), with the latter scaling differently with $K$. This assumption, along with the application of the central limit theorem following the approach in tree-like committee machines [7, 53], leads to a substantial simplification of the RS entropy and enables the treatment of general activation functions for large $K$, see the supplementary material (SM) for the complete derivation.
We note that the overlap structure assumed here is similar to that used in the capacity calculations for this architecture [46, 15] and in the teacher-student setting with spherical constraint [41] in the large-$K$ limit. Nevertheless, the absence of the spherical constraint, together with the introduction of the bias term in a non-Bayes optimal setting, distinguishes the present derivation. These additions are crucial to obtain well-defined RS equations for general activation functions and to explicitly analyze the specialization transition. Moreover, this setting is arguably relevant from a machine-learning perspective, as it corresponds to ERM rather than Bayesian learning.
@@
Given $(m^{*},q^{*},v^{*})$, a solution of the RS equations maximizing the RS potential (9), the RS approximation to the (Gibbs) generalisation error $\lim_{N\to\infty}\epsilon_{\rm Gibbs}$ is given by $\mathbb{E}_{(t,s)}[(f(t)-f(s))^{2}]$. Notice that $\mathcal{K}_{0}(d_{1},d_{2},a)$ coincides with the NNGP kernel [32, 52] of large-width neural networks at initialization. Similarly, the RS approximation to the training error $\lim_{N\to\infty}\epsilon_{\rm tr}$ reads $\partial_{\beta}\mathcal{G}_{E}|_{m^{*},q^{*},v^{*}}$.
Figure 1: Free entropy $\Phi_{\beta,K}^{\rm RS}$ (first panel) and order parameters associated with the stationary solutions of the RS equations, as a function of the sample-to-parameter ratio $\tilde{\alpha}$, for quadratic activation function $\sigma(x)=x^{2}$, MSE loss, weight decay $\lambda=0.1$, number of hidden units $K=10$ and $\beta=10$. Blue curves represent off-diagonal components of the teacher-student overlap $m^{*}$, the student-student overlap $q^{*}$, and the self-student overlap $q_{\rm self}^{*}$. Red curves represent the dominant diagonal components of these quantities. The dotted-dashed lines correspond to the permutation-symmetric branch (where $m_{d}^{*}=q_{d}^{*}=0$), the solid lines to the specialization branch ($m_{d}^{*}>0,q_{d}^{*}>0$) and the dashed one to the memorization branch ($m_{d}^{*}=0,q_{d}^{*}>0$).